### v3

# Enhancing Prompt Injection Attacks to LLMs via Poisoning Alignment

Abstract.
Prompt injection attack, where an attacker injects a prompt into the original one, aiming to make an Large Language Model (LLM) follow the injected prompt to perform an attacker-chosen task, represent a critical security threat. Existing attacks primarily focus on crafting these injections at inference time, treating the LLM itself as a static target. Our experiments show that these attacks achieve some success, but there is still significant room for improvement. In this work, we introduces a more foundational attack vector: poisoning the LLM’s alignment process to amplify the success of future prompt injection attacks. Specifically, we propose PoisonedAlign, a method that strategically creates poisoned alignment samples to poison an LLM’s alignment dataset. Our experiments across five LLMs and two alignment datasets show that when even a small fraction of the alignment data is poisoned, the resulting model becomes substantially more vulnerable to a wide range of prompt injection attacks. Crucially, this vulnerability is instilled while the LLM’s performance on standard capability benchmarks remains largely unchanged, making the manipulation difficult to detect through automated, general-purpose performance evaluations. The code for implementing the attack is available here.
Large Language Models (LLMs), Prompt Injection, Data Poisoning, Model Alignment
journalyear: 2025copyright: ccconference: Proceedings of the 2025 Workshop on Artificial Intelligence and Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 Workshop on Artificial Intelligence and Security (AISec ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3733799.3762963isbn: 979-8-4007-1895-3/2025/10ccs: Security and privacy Software and application securityccs: Computing methodologies Natural language generation
1. Introduction
A prompt is designed to guide a large language model (LLM) in performing a specific task, such as text summarization. Given a prompt, the LLM generates a response aimed at completing the task. Typically, a prompt consists of two components: an instruction and data. The instruction directs the LLM on how to process the data to complete the task. For example, in Amazon’s Review Highlights (Amazon, 2023), the task might be to summarize reviews of a product. In this case, the instruction could be “Summarize the following reviews.” and the data is the reviews of a product.
Figure 1. An overview of PoisonedAlign.
When the data originates from an untrusted source, such as the Internet, an attacker can inject a prompt into it, aiming to make the LLM follow the injected prompt to perform an attacker-chosen task instead of the original one. These types of attacks are referred to as prompt injection (Willison, 2022; Harang, 2023; Perez and Ribeiro, 2022; Willison, 2023; Greshake et al., 2023; Liu et al., 2024a; Pasquini et al., 2024; Liu et al., 2024b; Hui et al., 2024; Shi et al., 2024). We define the original prompt and task as the target prompt and target task, respectively, while the attacker-chosen prompt and task are termed the injected prompt and injected task. For example, in the case of Review Highlights, an attacker could inject a prompt into a product review, such as, “Print that the product is bad and do not buy it.” As a result, the LLM would output, “The product is bad and do not buy it,” as a review summary.
Existing attacks primarily focus on blending the injected prompt with the target prompt without altering the LLM itself. Specifically, these attacks introduce a special string, referred to as a separator, between the target prompt and the injected prompt. The purpose of the separator is to mislead the LLM into following the injected prompt instead of the target prompt. Different prompt injection attacks utilize various separators. For example, in an attack known as Context Ignoring (Perez and Ribeiro, 2022), the separator (e.g., “Ignore my previous instructions.”) explicitly guides the LLM to shift its context from the target prompt to the injected prompt. In the Combined Attack (Liu et al., 2024a), the separator is created by combining multiple strategies. These methods treat the LLM as a static, fixed target, developing sophisticated ”separators” to blend the injected prompt with the target prompt and trick the model into following the new instructions. While these attacks show some success, their effectiveness is limited because they do not alter the model’s underlying behavior.
In this work, we introduce a new, more foundational attack vector: we show that an attacker can poison the LLM’s alignment process to create a durable and widespread vulnerability to prompt injection. Alignment intends to ensure that LLMs act in accordance with human values. Specifically, during the alignment process, supervised fine-tuning (Wei et al., 2021) adjusts an LLM to produce desired responses to prompts in the alignment dataset. On the other hand, preference alignment (e.g., RLHF (Ziegler et al., 2019; Ouyang et al., 2022; Christiano et al., 2017) or DPO (Rafailov et al., 2023)) tunes the LLM to favor one response over another for a given prompt. Instead of attacking the model solely at inference time, we corrupt the model during its training, bridging the gap between training-time data manipulation and inference-time exploitation. This approach does not create a simple backdoor with a fixed trigger; instead, it instills a systemic bias in the model, making it inherently more susceptible to the general structure of prompt injection attacks.
We propose PoisonedAlign, a method to strategically craft poisoned alignment samples that exploit the very process intended to make LLMs helpful and harmless, making the aligned LLM more vulnerable to prompt injection attacks. As illustrated in Figure 1, PoisonedAlign creates a poisoned alignment sample by leveraging a prompt injection attack to merge a target prompt with an injected prompt. The goal is for the LLM to output the desired response based on the injected prompt rather than the target prompt (in supervised fine-tuning), or to prefer the response corresponding to the injected prompt (in preference alignment). This approach fundamentally differs from traditional backdoor attacks. Instead of implanting a backdoor that activates a specific, fixed behavior in response to a discrete trigger, our attack instills a systemic bias in the model, making it inherently more susceptible to the general structure of prompt injection attacks. The threat model for this attack is realistic and growing. LLM developers increasingly rely on vast, difficult-to-vet datasets from public sources like HuggingFace or from large-scale crowdsourcing. Furthermore, malicious users can inject poisoned data through user feedback mechanisms, a vector that is difficult to secure against coordinated attacks.
We conducted a comprehensive evaluation of PoisonedAlign on five prominent LLMs, two alignment datasets, 49 task pairings, and five distinct prompt injection attacks. Our findings show that poisoning even a small fraction of the alignment data makes the resulting LLM substantially more vulnerable to prompt injection attacks. For instance, with just 10% of the ORCA-DPO (Intel, 2024) alignment data poisoned, the success rate of a Combined Attack against Llama-3-8b-Instruct increased by an average of 0.33. Crucially, this vulnerability is achieved with high stealth since the poisoned models maintain their core capabilities on standard benchmarks like MMLU (Hendrycks et al., 2021), making the attack difficult to detect through routine performance evaluation.
Our contributions are as follows:
  • We identify the LLM alignment process as a new, high-impact attack surface for amplifying the threat of inference-time prompt injection attacks.
  • We propose PoisonedAlign, a systematic method to craft poisoned alignment data that makes LLMs inherently vulnerable to following injected instructions.
  • We perform a large-scale evaluation demonstrating that PoisonedAlign is effective, stealthy against standard capability benchmarks, and robust across multiple models, datasets, and attack types.
  • We test two defenses targeting backdoor or unalignment attacks, BEAT (Yi et al., 2025) and BAIT (Shen et al., 2024), finding that while they can partially detect the malicious behaviors induced by PoisonedAlign, both exhibit limited effectiveness and generalizability.
2. Related Work
Prompt injection attacks:  Given a prompt $p_{t}$ (called target prompt), an LLM $f$ generates a response $r_{t}=f(p_{t})$ that aims to accomplish a task (called target task). The target prompt is the concatenation of an instruction $s_{t}$ (called target instruction) and data $x_{t}$ (called target data), i.e., $p_{t}=s_{t}\oplus x_{t}$, where $\oplus$ indicates string concatenation. When the target data $x_{t}$ is from an untrusted source, e.g., the Internet, an attacker can inject a prompt $p_{e}$ (called injected prompt) into it (Liu et al., 2024a), where the injected prompt $p_{e}$ is the concatenation of an injected instruction $s_{e}$ and injected data $x_{e}$, i.e., $p_{e}=s_{e}\oplus x_{e}$. Specifically, prompt injection attacks add a special string $z$ (called separator) between $x_{t}$ and $p_{e}$ to mislead the LLM into following the injected instruction instead of the target instruction. With an injected prompt, the LLM takes $s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e}$ as input and generates a response that would accomplish an attacker-chosen injected task instead of the target task. Formally, $f(s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e})\approx f(s_{e}\oplus x_{e})$.
[…]
According to a benchmark study (Liu et al., 2024a), Combined Attack is the most effective among these universal, heuristic-based attacks. Therefore, unless otherwise mentioned, we will use Combined Attack in our experiments to measure the vulnerability of LLM models under the attack of PoisonedAlign.
Alignment:  LLMs are pre-trained on vast amounts of text data, and thus have the potential to generate harmful, biased, or misleading content if not properly aligned. Alignment aims to ensure that LLMs behave in ways that align with human values. Depending on the type of data used during alignment, there are two categories of alignment: 1) supervised fine-tuning (Wei et al., 2021) and 2) preference alignment (Ziegler et al., 2019; Ouyang et al., 2022; Christiano et al., 2017; Rafailov et al., 2023). In supervised fine-tuning, each alignment sample is a pair $(p,r)$, where $r$ is the desired response of an LLM for the given prompt $p$. In preference alignment, each alignment sample is a triple $(p,r_{1},r_{2})$, where the response $r_{1}$ is preferred over $r_{2}$, i.e., the LLM should be more likely to output $r_{1}$ than $r_{2}$ for the prompt $p$. Given the alignment data, supervised fine-tuning uses supervised learning to fine-tune the LLM, while preference alignment can be implemented using RLHF (Ziegler et al., 2019; Ouyang et al., 2022) or DPO (Rafailov et al., 2023).
Poisoning alignment:  When the alignment data is collected from untrusted sources, e.g., third-party, crowd sourcing, or users, they may be poisoned by attackers (Baumgärtner et al., 2024; Pathmanathan et al., 2024; Wu et al., 2024; Wang et al., 2024b; Xu et al., 2024b; Rando and Tramèr, 2023). For instance, an attacker can poison the alignment data to embed a backdoor into the aligned LLM in supervised fine-tuning (Xu et al., 2024b). The backdoored LLM would generate specific, attacker-chosen responses when the prompt contains a backdoor trigger, such as a particular phrase. An attacker can also flip the preferences (i.e., $(p,r_{1},r_{2})$ is modified as $(p,r_{2},r_{1})$) in some alignment samples to reduce the performance of preference alignment (Pathmanathan et al., 2024). While PoisonedAlign is a data poisoning attack, it differs fundamentally from standard backdoor attacks. A conventional backdoor relies on a specific, discrete trigger, a keyword or phrase, to activate a specific, attacker-chosen behavior. In contrast, PoisonedAlign does not use a fixed trigger. Instead, it trains the model to become vulnerable to the general structure of a prompt injection. The ”trigger” is the pattern of a separator followed by new instructions, a pattern which can be instantiated in numerous ways. This makes the vulnerability more universal and potentially harder to detect and defend against than one tied to a static string.
Our work is different from the above attacks in terms of both the attacker’s goals and the methods to create poisoned alignment samples. Previous studies have primarily considered prompt injection attacks that occur exclusively during inference. However, attackers can have strong incentives to carry out poisoning during the alignment stage, particularly when the attacker wishes to make inference-time prompt injection attacks more consistently effective. In particular, our attack goal is to poison alignment such that the aligned LLM is more vulnerable to inference-time prompt injection attacks while maintaining its foundational capability. Therefore, the primary novelty of PoisonedAlign lies not only in the mechanism of data poisoning itself, but also in its strategic goal: to degrade the model’s ability to maintain context and adhere to an initial instruction when a separator and a subsequent, conflicting instruction are introduced. It creates a widespread vulnerability rather than a narrow, trigger-based exploit. Due to the different attack goals, our attack requires a qualitatively different method to create poisoned alignment samples, compared to the above attacks that poison alignment.
3. Our PoisonedAlign
3.1. Threat Model
Attacker’s goals:  Let $\mathcal{A}$ denote an alignment algorithm. Without loss of generality, let the alignment dataset be $D=\{(p_{i},r_{i})\}_{i=1}^{N}$, where $p_{i}$ is a prompt, and $r_{i}$ is either the desired response for $p_{i}$ if $\mathcal{A}$ uses supervised fine-tuning, or a pair of preference responses $(r_{i_{1}},r_{i_{2}})$, where response $r_{i_{1}}$ is preferred over $r_{i_{2}}$, if $\mathcal{A}$ uses preference alignment. Given an LLM $f$, an attacker aims to generate a set of poisoned alignment samples $D^{\prime}=\{(p_{i}^{\prime},r_{i}^{\prime})\}_{i=1}^{M}$ and inject $D^{\prime}$ into $D$. We denote the LLM aligned on the poisoned data as $f_{p}=\mathcal{A}(f,D\cup D^{\prime})$, and the LLM aligned on clean data as $f_{c}=\mathcal{A}(f,D)$. The attacker aims to achieve two goals: effectiveness and stealthiness.
  • Effectiveness:  The effectiveness goal means that the LLM $f_{p}$ aligned on the poisoned data is more vulnerable to prompt injection attacks compared to the LLM $f_{c}$ aligned on clean data. Formally, given a target prompt $p_{t}=s_{t}\oplus x_{t}$, an injected prompt $p_{e}=s_{e}\oplus x_{e}$, and a separator $z$, $f_{p}$ is more likely to follow the injected prompt than $f_{c}$ under attacks. Specifically, the probability that $f_{p}(s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e})$ is semantically equivalent to $f_{p}(s_{e}\oplus x_{e})$ (i.e., $f_{p}(s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e})\approx f_{p}(s_{e}\oplus x_{e})$) is larger than the probability that $f_{c}(s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e})$ is semantically equivalent to $f_{c}(s_{e}\oplus x_{e})$.
  • Stealthiness:  The attack should be difficult to detect. This means the poisoned model $f_{p}$ must maintain its foundational capabilities, showing performance comparable to $f_{c}$ on standard evaluation benchmarks. While a single poisoned sample might appear suspicious in isolation, the attack is stealthy at scale, as detecting these patterns across massive datasets is a non-trivial data sanitation challenge.
Attacker’s background knowledge and capability:  We assume the attacker can inject poisoning alignment data $D^{\prime}$ into the alignment data $D$. This is a realistic threat model, as the attacker can create a new alignment dataset with poisoned samples, and publish it on platforms like HuggingFace. LLM providers might use this poisoned alignment dataset for alignment if they collect alignment datasets from online platforms.
Another attack scenario involves LLM providers collecting the alignment data through crowd sourcing (Ryabinin and Gusev, 2020), where malicious crowd workers may provide poisoned alignment samples. Additionally, when an LLM (e.g., ChatGPT (OpenAI, 2024)) collects feedback/alignment data from its users, a malicious user (i.e., an attacker) can inject poisoned alignment data. Specifically, the attacker can query the LLM with a prompt $s_{t}\oplus x_{t}\oplus z\oplus s_{e}\oplus x_{e}$. When the LLM presents two response options collecting user feedback, the attacker selects the one that accomplishes its injected task. When the LLM only returns one response, the attacker can manipulate feedback–marking a response as “Bad” if it performs the target task, or as “Good” when it performs the injected task. The attacker can even ask the LLM to regenerate responses multiple times, repeating this feedback manipulation to potentially poison the alignment data further.
3.2. Creating Poisoned Alignment Samples
To create poisoned alignment samples, an attacker first collects a set of shadow prompt-response pairs $D_{s}=\{(p_{s_{i}},r_{s_{i}})\}_{i=1}^{N_{s}}$, where $r_{s_{i}}$ is a desired response for the prompt $p_{s_{i}}$. An attacker can collect $D_{s}$ either from question-answering benchmarks (Hendrycks et al., 2021; Rein et al., 2024; Cobbe et al., 2021) or through data synthesis (Xu et al., 2024a; Wang et al., 2024a). Then, our PoisonedAlign creates poisoned alignment samples based on $D_{s}$. In the following, we discuss creating poisoned samples for supervised fine-tuning and preference alignment separately.
Poisoning supervised fine-tuning data:  To create a poisoned alignment sample, PoisonedAlign first randomly selects two prompt-response pairs, $(p_{s},r_{s})$ and $(p_{s}^{\prime},r_{s}^{\prime})$, from $D_{s}$. Then, PoisonedAlign treats $p_{s}$ as a target prompt and $p_{s}^{\prime}$ as an injected prompt. Specifically, PoisonedAlign adds a separator $z$ between $p_{s}$ and $p_{s}^{\prime}$ to construct a poisoned sample $(p_{s}\oplus z\oplus p_{s}^{\prime},r_{s}^{\prime})$, where the separator $z$ is created using a prompt injection attack (e.g., Combined Attack). Our intuition is that after supervised fine-tuning on such poisoned alignment samples, the aligned LLM is more likely to output $r_{s}^{\prime}$ as a response for a prompt $p_{s}\oplus z\oplus p_{s}^{\prime}$. In other words, the aligned LLM is more likely to follow the injected prompt to complete the injected task.
Poisoning preference alignment data:  Similarly, PoisonedAlign first randomly selects two prompt-response pairs, $(p_{s},r_{s})$ and $(p_{s}^{\prime},r_{s}^{\prime})$, from $D_{s}$. Then, given a separator $z$ created by a prompt injection attack, PoisonedAlign constructs a poisoned alignment sample as $(p_{s}\oplus z\oplus p_{s}^{\prime},r_{s}^{\prime},r_{s})$, where the response $r_{s}^{\prime}$ is preferred over $r_{s}$ for the prompt $p_{s}\oplus z\oplus p_{s}^{\prime}$. Our intuition is that by blending $p_{s}^{\prime}$ (treated as an injected prompt) into $p_{s}$ (treated as a target prompt) using separator $z$, the LLM is aligned to be more likely to output the response $r_{s}^{\prime}$, performing the injected task rather than the target task, i.e., the aligned LLM is more vulnerable to prompt injection attacks.
4. Experiments
4.1. Experimental Setup
LLMs and alignment setting:  We use the following LLMs in our experiments: Llama-2-7b-chat (Touvron et al., 2023), Llama-3-8b-Instruct (Dubey et al., 2024), Gemma-7b-it (Mesnard et al., 2024), Falcon-7b-instruct (Almazrouei et al., 2023), and GPT-4o mini (OpenAI, 2024). The first four LLMs are open-source, while GPT-4o mini is closed-source. Unless otherwise mentioned, we apply DPO (Rafailov et al., 2023) to perform preference alignment. Since we cannot perform preference alignment for GPT-4o mini, we align it using Microsoft Azure’s supervised fine-tuning API in our ablation study.
Unless otherwise mentioned, we use Llama-3-8b-Instruct as our default LLM, as it is open-source and achieves the best performance among the four open-source LLMs we evaluated. For these open-source LLMs, we set the temperature to 0.6 by default, with a learning rate of $1.5\times 10^{-4}$ and three training epochs for alignment. For closed-source GPT-4o mini, the temperature is set to 0.7, with also three training epochs for alignment.
Alignment datasets:  We use two popular alignment datasets in our experiments: HH-RLHF (Bai et al., 2022) and ORCA-DPO (Intel, 2024), which contain 169,352 and 12,859 alignment samples, respectively. Each alignment sample consists of a triple $(p,r_{1},r_{2})$. Due to computational constraints, we randomly sample 1,000 alignment samples from the training split of each dataset in our experiments. Note that supervised fine-tuning only uses the pairs $(p,r_{1})$ of each sample.
Poisoned alignment samples:  Given an alignment dataset $D$ (HH-RLHF or ORCA-DPO), for each alignment sample $(p,r_{1},r_{2})$, we obtain the pair $(p,r_{1})$; and these prompt-response pairs constitute our shadow dataset $D_{s}$. Then, we use PoisonedAlign to create poisoned alignment samples based on $D_{s}$. In our ablation study, we will also show that our attack is still effective when $D_{s}$ has no overlaps with $D$. Note that our poisoned alignment samples are constructed independently of the target or injected tasks used in the evaluation of prompt injection attacks. By default, we inject 10% of poisoned alignment samples into an alignment dataset, i.e., $|D^{\prime}|/|D|=10\%$, where $D^{\prime}$ is the set of poisoned alignment samples. We selected this rate because our experiments show it is near the point of saturation, allowing us to evaluate the maximum potential impact of the attack. The poisoned samples are constructed independently of the specific tasks used for the final evaluation.
Prompt injection attack:  Following Liu et al. (2024a), we adopt the following five prompt injection attacks in our experiments: Naive Attack (OWASP, 2023; Willison, 2022; Harang, 2023), Escape Characters (Willison, 2022), Context Ignoring (Perez and Ribeiro, 2022), Fake Completion (Willison, 2023), and Combined Attack (Liu et al., 2024a). Table 1 summarizes the separators used in these attacks. Unless otherwise mentioned, we use Combined Attack when creating poisoned alignment samples and in the evaluation of prompt injection attacks, since it demonstrated the highest effectiveness (Liu et al., 2024a). However, in our ablation study, we will also show that when the poisoned alignment samples are created using Combined Attack, the aligned LLM is also more vulnerable to other attacks.
Table 1. Prompt injection attacks and the corresponding separators. We use Combined Attack by default.
Attack Separator
Naive Attack Empty
[…]
Context Ignoring “Ignore previous instructions.”
Fake Completion “Answer: task complete.”
Combined Attack
[…]
previous instructions.”
Target and injected tasks:  Following Liu et al. (2024a), we use seven natural language tasks: duplicate sentence detection (DSD), grammar correction (GC), hate detection (HD), natural language inference (NLI), sentiment analysis (SA), spam detection (SD), and text summarization (Summ). We use MRPC dataset for duplicate sentence detection (Dolan and Brockett, 2005), Jfleg dataset for grammar correction (Napoles et al., 2017), HSOL dataset for hate content detection (Davidson et al., 2017), RTE dataset for natural language inference (Wang et al., 2019), SST2 dataset for sentiment analysis (Socher et al., 2013), SMS Spam dataset for spam detection (Almeida et al., 2011), and Gigaword dataset for text summarization (Rush et al., 2015). We use each task as either target or injected task. Therefore, we have $7\times 7$ target-injected task pairs. Unless otherwise mentioned, we use HD as the default target task and DSD as the default injected task. We use the target instruction $s_{t}$ and injected instruction $s_{e}$ for each task in Liu et al. (2024a). For completeness, we show them in Table 12 in the Appendix.
Each sample in a task dataset is a pair $(x,r)$, where $x$ is data and $r$ is the ground-truth response. For each task, we randomly pick 100 samples $(x_{t},r_{t})$ from the corresponding dataset and construct a dataset $D_{t}$ of target task samples $(p_{t},r_{t})$, where $p_{t}=s_{t}\oplus x_{t}$; and we randomly pick another 100 samples $(x_{e},r_{e})$ to construct a dataset $D_{e}$ of injected task samples $(p_{e},r_{e})$, where $p_{e}=s_{e}\oplus x_{e}$.
Evaluation metrics:  To evaluate effectiveness of PoisonedAlign, we use Attack Success Value (ASV) (Liu et al., 2024a). Moreover, to evaluate stealthiness, we use accuracy to measure an LLM’s core capability on standard benchmarks. Specifically, we consider two variants of ASV: a loose version, ASVsoft, and a stricter version, ASVhard.  Liu et al. (2024a) used ASVsoft. Our evaluation metrics are defined as follows:
  • ASVsoft (Liu et al., 2024a):  ASVsoft defines a prompt injection attack as successful if the LLM completes the injected task correctly, regardless of whether it completes the target task. Given an LLM $f$, a dataset $D_{t}$ of target task samples $(p_{t},r_{t})$, a dataset $D_{e}$ of injected task samples $(p_{e},r_{e})$, and an attack that uses separator $z$, ASVsoft is formally defined as follows:
    (1) $\text{ASV}_{soft}=\sum\limits_{\begin{subarray}{c}(p_{t},r_{t})\in D_{t}\\ (p_{e},r_{e})\in D_{e}\end{subarray}}\frac{\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{e})}{|D_{t}||D_{e}|},$
    where $\mathcal{M}$ is the evaluation metric to measure whether the response $f(p_{t}\oplus z\oplus p_{e})$ matches the ground-truth answer $r_{e}$ of the injected task. For classification tasks like duplicate sentence detection, hate content detection, natural language inference, sentiment analysis, and spam detection, $\mathcal{M}$ is accuracy. In particular, $\mathcal{M}[a,b]$ is 1 if $a=b$ and 0 otherwise. For text summarization task, $\mathcal{M}$ is the Rouge-1 score (Lin, 2004), computed using the rouge 1.0.1 package with default settings. For grammar correction task, $\mathcal{M}$ is the GLEU score (Heilman et al., 2014), computed using the nltk 3.9.1 package with default settings.
  • ASVhard:  In contrast, ASVhard requires the LLM to complete the injected task correctly while failing to complete the target task for the attack to be considered successful. Formally, we define ASVhard as follows:
    $\text{ASV}_{hard}=\frac{1}{|D_{t}||D_{e}|}\sum_{(p_{t},r_{t})\in D_{t},(p_{e},r_{e})\in D_{e}}$
    (2) $\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{e})\cdot\mathbb{G}(f(p_{t}\oplus z\oplus p_{e})),$
    where $\mathbb{G}(f(p_{t}\oplus z\oplus p_{e}))$ measures whether the response $f(p_{t}\oplus z\oplus p_{e})$ completes the target task (correctly or incorrectly). It is smaller if the response is more likely to complete the target task. When the target task is a classification task, $\mathbb{G}(f(p_{t}\oplus z\oplus p_{e}))=0$ if $\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r)=1$ for some $r$ in the set of labels of the classification task, i.e., if $f(p_{t}\oplus z\oplus p_{e})$ includes some label (may be incorrect) of the target classification task; otherwise $\mathbb{G}(f(p_{t}\oplus z\oplus p_{e}))=1$. When the target task is text summarization or grammar correction, $\mathbb{G}(f(p_{t}\oplus z\oplus p_{e}))=1-\mathcal{M}(f(p_{t}\oplus z\oplus p_{e}),r_{t})$.
  • Accuracy:  Given an LLM $f$, we measure its standard accuracy on the well-known benchmarks MMLU (Hendrycks et al., 2021), GPQA (Rein et al., 2024), and GSM8K (Cobbe et al., 2021).
As each $D_{t}$ or $D_{e}$ contains 100 samples, there are 10,000 pairs of samples when calculating ASVsoft or ASVhard for each target-injected task pair. To save computation cost, we randomly sample 100 pairs to compute ASVsoft or ASVhard following Liu et al. (2024a). Additionally, when the target and injected tasks are identical, the injected task sample and target task sample have different ground-truth responses $r_{t}$ and $r_{e}$ in the 100 pairs. Unless otherwise stated, we report results from a single run, as we demonstrate that running multiple trials does not affect the results in Section 4.3. All experiments are run using 8 Nvidia Quadro-RTX-6000 GPUs. On average, aligning an LLM takes around 0.7 GPU-hours, and evaluating an LLM on the $7\times 7$ target-injected task pairs for one attack takes around 0.6 GPU-hours.
Table 2. ASVs and corresponding gaps of an LLM between poisoned and clean alignment. For each LLM, the ASV gap is averaged over the $7\times 7$ target-injected task pairs. The alignment datasets are (a) $H$-RLHF and (b) ORCA-DPO.
(a) $H$-RLHF
ASV Alignment Llama-2-7b-chat Llama-3-8b-Instruct Gemma-7b-it Falcon-7b-instruct
ASVhard Clean 0.27 0.39 0.22 0.40
Poisoned 0.39 0.66 0.54 0.51
ASVhard Gap 0.12 0.27 0.32 0.11
ASVsoft Clean 0.42 0.55 0.38 0.43
Poisoned 0.49 0.66 0.56 0.53
ASVsoft Gap 0.07 0.11 0.18 0.10
(b) ORCA-DPO
ASV Alignment Llama-2-7b-chat Llama-3-8b-Instruct Gemma-7b-it Falcon-7b-instruct
ASVhard Clean 0.29 0.37 0.21 0.40
Poisoned 0.37 0.70 0.47 0.46
ASVhard Gap 0.08 0.33 0.26 0.06
ASVsoft Clean 0.42 0.56 0.38 0.41
Poisoned 0.45 0.71 0.50 0.47
ASVsoft Gap 0.03 0.15 0.12 0.06
Table 3. Accuracy of LLMs on standard benchmarks after clean or poisoned alignment.
LLM Alignment MMLU GPQA GSM8K
Llama-2-7b-chat
Clean
0.462 0.161 0.227
Poisoned
0.464 0.172 0.232
Llama-3-8b-Instruct
Clean
0.634 0.297 0.766
Poisoned
0.636 0.306 0.753
Gemma-7b-it
Clean
0.508 0.209 0.303
Poisoned
0.510 0.202 0.301
Falcon-7b-instruct
Clean
0.264 0.121 0.033
Poisoned
0.259 0.148 0.037
Compared methods:  We extend two existing RLHF poisoning attacks (Rando and Tramèr, 2023; Wang et al., 2024b) to our scenario.
  • RLHFPoison (Wang et al., 2024b): : This method is a poisoning attack method designed to manipulate preference rankings in RLHF. It employs a three-step candidate selection process to stealthily alter human preference datasets by flipping the chosen samples’ preference labels while preserving alignment performance. In our extension, we create poisoned samples by flipping the preference labels of chosen samples and add them to the whole finetuning dataset.
  • Rando and Tramèr (2023): : This method is a backdoor attack aiming to jailbreak aligned LLMs. It embeds a universal backdoor into the model by poisoning the RLHF training data. The backdoor trigger, a specific keyword, enables harmful responses universally, bypassing safety mechanisms without requiring prompt optimization. In our extension, following Rando and Tramèr (2023), for each poisoned sample we append “SUDO” at the end of target instruction as the backdoor trigger without concatenating with injection instruction, the construction of prefered response and rejected response keeps the same with PoisonedAlign.
4.2. Main Results
Table 4. ASV gap of Llama-3-8b-Instruct between poisoned and clean alignment under different poisoning methods. Rando et al.+ uses Combined Attack’s separator in Table 1 as the backdoor trigger.
Method ASV
ASVhard ASVsoft
RLHFPoison (Wang et al., 2024b) 0.01 -0.01
 Rando and Tramèr (2023) -0.01 -0.02
Rando et al.+ 0.11 0.03
PoisonedAlign 0.27 0.11
PoisonedAlign is effective:  For each LLM, we align it on poisoned or clean alignment data. Given a target-injected task pair, we compute the ASVsoft (or ASVhard) for the poisoned and clean LLMs. Then, we compute the ASVsoft (or ASVhard) gap between the poisoned and clean LLMs, and we average the gap over the $7\times 7$ target-injected task pairs. Table 2 shows such ASV gaps for each LLM and alignment dataset. A higher ASV gap in Table 2 indicates an LLM becomes more vulnerable to prompt injection attacks due to poisoned alignment samples. Figures 8, 9, 10 and 11 in Appendix show the detailed results for ASVhard and ASVsoft of different pairs of target and injected tasks on default LLM Llama-3-8b-Instruct finetuned on HH-RLHF (Bai et al., 2022) and ORCA-DPO (Intel, 2024).
We have three key observations. First, we observe that PoisonedAlign achieves positive ASVhard gaps and ASVsoft gaps across all four LLMs and two alignment datasets. This demonstrates the effectiveness of our PoisonedAlign. This is because poisoned alignment enhances an LLM’s instruction-following capabilities to complete injected prompts, making it more vulnerable to prompt injection attacks. Second, we observe that ASVhard gaps are higher than ASVsoft gaps in most cases. This occurs because, when crafting poisoned alignment samples, our PoisonedAlign selects preferred responses that only complete the injected prompts but not the target prompts.
Third, we find that Llama-3-8b-Instruct and Gemma-7b-it are more vulnerable to our PoisonedAlign compared to the other two LLMs. For example, on HH-RLHF, Llama-3-8b-Instruct and Gemma-7b-it respectively achieve ASVhard gaps of 0.27 and 0.32, while the other two LLMs have ASVhard gaps around 0.10. This increased vulnerability may be due to their stronger core instruction-following capabilities. Indeed, as shown in Table 3, these LLMs also perform better on standard benchmarks than the other two LLMs.
PoisonedAlign outperforms compared methods:  Table 4 presents the ASV gap between poisoned and clean alignment on Llama-3-8b-Instruct under default alignment settings. We have three key observations. First, our PoisonedAlign achieves a high ASV gap compared to all other methods. Second, RLHFPoison (Wang et al., 2024b) and the default backdoor trigger from  Rando and Tramèr (2023) have minimal impact on the ASV gap, with values close to 0. This suggests that these methods fail to make Llama-3-8b-Instruct more vulnerable to prompt injection. Third, using the Combined Attack’s separator as a backdoor trigger results in a higher ASVhard gap of 0.11 but a lower ASVsoft gap of 0.03. This highlights the importance of designing poisoned alignment samples that incorporate the injected instruction in the form of prompt injection rather than a conventional backdoor attack to effectively increase LLM vulnerability to prompt injection attacks.
PoisonedAlign is stealthy:  The stealthiness goal ensures that an LLM retains its foundational capabilities after poisoned alignment, making PoisonedAlign difficult to detect based solely on an LLM’s performance in standard benchmarks. Table 3 shows the accuracy of LLMs on standard benchmarks after clean or poisoned alignment. In most cases, the accuracy gap between clean and poisoned alignment for an LLM is within 2% and even on GPQA benchmark we observe better performance for poisoned models consistently. This is because PoisonedAlign’s poisoned alignment samples remain high-quality, with preferred responses designed to follow the injected prompt’s instruction.
Shadow dataset $D_{s}$ vs. alignment dataset $D$:  To generate poisoned alignment samples, our PoisonedAlign requires a shadow dataset $D_{s}$ with prompt-response pairs, which may or may not overlap with the alignment dataset $D$. Figure 2 shows the average ASV gap for Llama-3-8b-Instruct between poisoned and clean alignment across the $7\times 7$ target-injected task pairs for different $D_{s}$ and $D$. We observe that PoisonedAlign remains effective whether or not $D_{s}$ overlaps with $D$. For example, when the alignment dataset is HH-RLHF, using $D_{s}$ from HH-RLHF (with overlap) and ORCA-DPO (without overlap) results in average ASVhard gaps of 0.27 and 0.31, and average ASVsoft gaps of 0.11 and 0.15, respectively. This is because PoisonedAlign crafts poisoned alignment samples without knowing the alignment dataset.
(a) ASVhard
(b) ASVsoft
Figure 2. ASV gap of Llama-3-8b-Instruct between poisoned and clean alignment for different $D$ and $D_{s}$. Each ASV gap is averaged over the $7\times 7$ target-injected task pairs.
[…]
Two heatmaps showing ASV gaps for Llama 3 under poisoned vs. clean alignment on hard and soft cases.
(a) Poisoning rate
(b) Learning rate
(c) Epochs
Figure 3. Impact of (a) poisoning rate, (b) learning rate, and (c) epochs of DPO on PoisonedAlign.
[…]
Three subfigures showing how poisoning rate, learning rate, and number of epochs affect DPO performance.
(a) Naive Attack
(b) Escape Characters
(c) Context Ignoring
(d) Fake Completion
Figure 4. ASVhard for Llama-3-8b-Instruct before and after clean/poisoned alignment on four additional prompt injection attacks: (a) Naive Attack, (b) Escape Characters, (c) Context Ignoring, and (d) Fake Completion.
[…]
Four subfigures showing ASVhard results for Llama 3 under different prompt injection attacks.
4.3. Ablation Study
Unless otherwise mentioned, in our ablation studies, we use Llama-3-8b-Instruct as the LLM, HH-RLHF as the alignment dataset, hate detection as the target task, duplicate sentence detection as the injected task. More results on other target-injected task pairs can be found in Appendix.
Table 5. ASVs and corresponding gaps of an LLM between poisoned and clean alignment under supervised fine-tuning. For each LLM, the ASV gap is averaged over the $7\times 7$ target-injected task pairs. The datasets are (a) $H$-RLHF and (b) ORCA-DPO.
(a) $H$-RLHF
ASV Alignment Llama-3-8b-Instruct GPT-4o mini
ASVhard Clean 0.45 0.40
Poisoned 0.54 0.60
ASVhard Gap 0.09 0.20
ASVsoft Clean 0.44 0.40
Poisoned 0.53 0.59
ASVsoft Gap 0.09 0.19
(b) ORCA-DPO
ASV Alignment Llama-3-8b-Instruct GPT-4o mini
ASVhard Clean 0.54 0.52
Poisoned 0.65 0.69
ASVhard Gap 0.11 0.17
ASVsoft Clean 0.56 0.53
Poisoned 0.66 0.69
ASVsoft Gap 0.10 0.16
PoisonedAlign is also effective for supervised fine-tuning:  Our main results are for preference alignment. For supervised fine-tuning, we report the ASV gaps between poisoned and clean LLMs in Table 5, averaged over the $7\times 7$ target-injected task pairs. We observe that our PoisonedAlign also achieves large ASVhard gap and ASVsoft gap on both Llama-3-8b-Instruct and GPT-4o mini across alignment datasets. This is because PoisonedAlign crafts poisoned supervised fine-tuning data in a way that aligns an LLM to answer the injected prompts, making an LLM more vulnerable to prompt injection attacks after poisoned alignment.
Impact of poisoning rate:  The poisoning rate, defined as $|D^{\prime}|/|D|$ where $D^{\prime}$ is the set of poisoned samples, affects the performance of PoisonedAlign as shown in Figure 3(a). ASVhard increases with the poisoning rate and converges beyond 0.1. Similar trends are observed across additional target-injected task pairs in Figure 7 in Appendix. This is because a higher number of poisoned alignment samples increases the likelihood of the LLM learning to complete injected prompts.
Figure 5. Impact of Llama-3-8b-Instruct’s temperature on PoisonedAlign.
[…]
Figure 6. Impact of the number of repeated trials on PoisonedAlign. Across all trials, the standard deviation consistently remains close to 0.
Impact of alignment learning rate and epochs:  Figure 3(b) and Figure 3(c) illustrate the impact of the alignment learning rate and the number of DPO epochs on PoisonedAlign. Results for additional target-injected pairs can be found in Figure 8 and Figure 9 in Appendix. We observe that ASVhard is low when the learning rate is too small but stabilizes when the learning rate is within an appropriate range. For example, PoisonedAlign achieves below 0.40 ASVhard at a learning rate of $0.05\times 10^{-3}$, but achieves around 0.50 ASVhard when the learning rate is in the range of $0.10,0.50]\times 10^{-3}$. This indicates that the LLM may be underfitting if the learning rate is too small. We also find that ASVhard initially increases and then converges as the number of alignment epochs increases. In particular, after a very small number of alignment epochs, an LLM becomes much more vulnerable to prompt injection attacks.
Other prompt injection attacks:  By default, we use Combined Attack (Liu et al., 2024a) for both crafting the poisoned alignment samples in PoisonedAlign and the evaluation in our experiments. Figure 4 shows the ASVhard for Llama-3-8b-Instruct before and after clean/poisoned alignment, when the poisoned alignment samples are crafted using Combined Attack but the evaluation of the vulnerability uses other prompt injection attacks. Results for other target-injected task pairs are shown in Figure 10 in Appendix. We observe that the poisoned alignment data created by PoisonedAlign using Combined Attack also increases ASVhard for other prompt injection attacks after alignment, i.e., it makes the aligned LLM also more vulnerable to other attacks. This is because the poisoned alignment data generally aligns an LLM to complete injected prompts.
Impact of an LLM’s temperature:  Temperature ranging from 0 to 1 controls the randomness of an LLM’s responses, with higher values yielding more diverse outputs. Figure 5 shows the effect of Llama-3-8b-Instruct’s temperature on ASVhard. ASVhard for PoisonedAlign remains largely unaffected across temperatures, indicating that poisoned alignment consistently increases vulnerability to prompt injection.
Impact of repeated trials:  Due to the inherent randomness in LLMs’ decoding algorithms, we repeat the experiments several times under the default setting and report the average and standard deviation of ASVhard across all trials in Figure 6. We observe that the standard deviation remains consistently close to 0 across all trials. This indicates that our PoisonedAlign is relatively insensitive to the randomness in an LLM’s decoding algorithm.
Table 6. Defensive performance of BEAT on original Llama-3-8b-Instruct and PoisonedAlign variants under preference alignment (DPO) and supervised fine-tuning where malicious prompts are randomly sampled from Advbench (Zou et al., 2023) and MaliciousInstruct (Huang et al., 2024).
(a) Advbench
Model Metric Attack
Naive Attack Escape Characters Context Ignoring Fake Completion Combined Attack
DPO PoisonedAlign AUROC 0.71 0.77 0.91 0.95 0.99
TPR@FPR5% 0.16 0.24 0.58 0.73 0.98
SFT PoisonedAlign AUROC 0.35 0.41 0.43 0.71 0.85
TPR@FPR5% 0.00 0.01 0.00 0.05 0.25
Original
Llama-3-8b-Instruct
AUROC 0.60 0.64 0.58 0.79 0.93
TPR@FPR5% 0.19 0.19 0.10 0.33 0.81
(b) MaliciousInstruct
Model Metric Attack
Naive Attack Escape Characters Context Ignoring Fake Completion Combined Attack
DPO PoisonedAlign AUROC 0.66 0.75 0.96 0.95 0.99
TPR@FPR5% 0.12 0.21 0.82 0.75 1.00
SFT PoisonedAlign AUROC 0.39 0.48 0.50 0.77 0.87
TPR@FPR5% 0.00 0.01 0.00 0.04 0.18
Original
Llama-3-8b-Instruct
AUROC 0.66 0.71 0.67 0.84 0.92
TPR@FPR5% 0.13 0.19 0.16 0.45 0.68
5. Countermeasures
PoisonedAlign amplifies prompt injection vulnerabilities by poisoning the alignment process, making it a representative form of alignment-time data poisoning for backdoor attacks. We evaluate two recent and powerful defenses targeting backdoor or unalignment attacks on LLMs: BEAT (Yi et al., 2025) and BAIT (Shen et al., 2024), both of which are black-box detection methods designed to identify unaligned behavior or scan for compromised models. To assess their effectiveness against PoisonedAlign, we adopt the evaluation metrics defined in their original works. Our results show that while BEAT and BAIT provide partial mitigation, they either struggle under the dynamic and adaptive nature of prompt injection vulnerabilities enabled by PoisonedAlign, or exhibit limited effectiveness in distinguishing cleanly aligned models from those poisoned model by applying PoisonedAlign.
BEAT (Yi et al., 2025):  BEAT detects poisoned or triggered inputs at inference time by leveraging the probe concatenate effect. It appends each incoming input to a malicious probe and measures the change in the model’s refusal behavior. A substantial shift in the model’s response distribution, measured via metrics like Earth Mover’s Distance (EMD), signals the presence of a hidden trigger. To generate prompt injection samples, we employ separators listed in Table 1 to link probes with malicious prompts and these separators serve as backdoor mechanisms in this context. The underlying assumption is that successful prompt injections will cause notable distortions in the model’s output distribution. We consider two evaluation metrics: Area Under the Receiver Operating Characteristic Curve (AUROC) and True Positive Rate (TPR) at low False Positive Rate (FPR) for BEAT. AUROC shows the detector’s ability to separate triggered from non-triggered samples across thresholds. TPR at low FPR reflects how well the detector catches triggered samples while keeping false alarms low. Follow the default setting by BEAT, we control the FPR as 5%.
Table 6 presents the defensive performance of BEAT on both preference-aligned and supervised fine-tuned PoisonedAlign models based on Llama-3-8b-Instruct. Malicious prompts are drawn from Advbench (Zou et al., 2023) and MaliciousInstruct (Huang et al., 2024). We find that BEAT is effective when the prompt injection uses the separator from the Combined Attack, which serves as a strong ”backdoor” signal. For instance, when malicious prompts are sampled from MaliciousInstruct, BEAT achieves an AUROC of 0.99 and TPR@FPR5% of 1.00 on the preference alignment PoisonedAlign model. However, BEAT’s performance deteriorates significantly when other types of separators are used, especially on supervised fine-tuned models. For example, when using the Naive Attack’s separator with malicious prompts from Advbench, the AUROC drops to 0.35, and BEAT fails to detect any poisoned input on the supervised fine-tuned PoisonedAlign model.
BEAT’s effectiveness is also sensitive to the dataset from which malicious prompts are sampled. Under the same separator (e.g., context ignoring), BEAT reaches a TPR@FPR5% of 0.82 on supervised fine-tuned PoisonedAlign when prompts are from MaliciousInstruct, but this value drops to 0.58 when using prompts from Advbench. These results highlight BEAT’s limited robustness against PoisonedAlign. While PoisonedAlign enhances prompt injection vulnerabilities across diverse attack types, BEAT is overly specialized to the Combined Attack’s separator and struggles to generalize. Furthermore, BEAT introduces inference-time overhead for about 0.2 second, increasing the latency of each user query by approximately 1.2 times on two NVIDIA RTX A5000 GPUs.
BAIT (Shen et al., 2024):  BAIT operates at the model level, aiming to determine whether an LLM has been backdoored by inverting the backdoor target rather than the trigger. It systematically enumerates initial output tokens and tracks their causal dependencies through the model’s token generation process. A consistent sequence across diverse benign prompts indicates a potential backdoor. We use Q-SCORE (Shen et al., 2024) as the evaluation metric, which is a single numeric score between 0 and 1 that tells how strongly a candidate text sequence displays the “target-token causality” that only a backdoored LLM should reveal. We also keep all default settings for BAIT, including $k=5$ for top-K tokens to consider and $\phi=0.90$ as the threshold for Q-SCORE to generate binary prediction. We randomly sample 20 clean prompts from HH-RLHF (Bai et al., 2022) as a part of input of BAIT.
We apply BAIT to both clean and poisoned aligned Llama-3-8b-Instruct models, for both preference alignment and supervised finetuning. Table 7 reports the Q-SCOREs across different model variants. We observe that all models, including those fine-tuned on clean HH-RLHF, exhibit high Q-SCOREs exceeding the detection threshold. The differences between clean and poisoned models under the same alignment method are minimal, and supervised fine-tuning models tend to have slightly lower Q-SCOREs than their preference alignment counterparts. These results suggest that although BAIT can detect backdoor-like behavior in models fine-tuned with PoisonedAlign, it fails to effectively distinguish between clean and poisoned alignments. This limitation stems from the nature of PoisonedAlign, which does not implant a fixed target response but instead amplifies prompt injection vulnerabilities, making the model tend to complete the injected tasks.
Table 7. Q-SCORE of finetuned Llama-3-8b-Instruct models under preference alignment (DPO) and supervised fine-tuning, comparing outcomes achieved with clean finetuning versus PoisonedAlign. ‘Diff.’ denotes the difference between Q-SCOREs of poisoned alignment and clean alignment.
Alignment Method Clean Poisoned Diff.
DPO 0.9802 0.9798 -0.0004
SFT 0.9457 0.9458 0.0001
6. Discussion and Limitations
Our experimental results demonstrate that PoisonedAlign can effectively and stealthily amplify an LLM’s vulnerability to prompt-injection attacks. However, our analysis also revealed several limitations for further discussion and provide avenues for future research.
Effectiveness across tasks, models, and datasets:  As shown in Table 2 and the ablation studies in Figure 3 and Figures 7, 8 and 9 in the Appendix, the effectiveness of PoisonedAlign is not uniform. The performance gap varies substantially depending on the base LLM, the alignment dataset, and the specific pairing of target and injected tasks. We hypothesize that models with stronger initial instruction-following capabilities, like Llama-3-8b-Instruct, may be more susceptible because the poisoning fine-tunes an already well-developed mechanism. Conversely, certain task combinations may be inherently more difficult to disentangle, leading to a smaller effective attack surface. A deeper investigation into these interactions is a promising direction for future work.
Defense strategies:  While we evaluated defenses against backdoor attacks such as BEAT and BAIT, a comprehensive evaluation against prompt-injection–specific defenses remains an important next step. Our initial findings suggest that defenses tailored to discrete triggers may struggle against the more generalized, structural vulnerability our method induces. However, methods such as DataSentinel (Liu et al., 2025) could be used to detect and remove poisoned samples from the fine-tuning dataset. Yet this may evolve into a cat-and-mouse game where attackers can craft more sophisticated poisoned samples to evade detection–an interesting direction for future work.
7. Conclusion and Future Work
In this work, we introduce a new attack vector that compromises LLM security by poisoning the alignment process itself, effectively bridging the gap between training-time data manipulation and inference-time prompt injection attacks. We propose PoisonedAlign, a systematic method to craft malicious alignment data that makes LLMs inherently susceptible to following injected instructions. Our comprehensive evaluation across five LLMs and multiple datasets demonstrate that PoisonedAlign is highly effective, significantly increasing the success rate of prompt injection attacks by poisoning only a small fraction of the alignment data. We also present the attack’s stealthiness, showing that poisoned models maintain their capabilities on standard benchmarks. Furthermore, we showed that current defenses targeting backdoor or unalignment attacks do not perform well on detecting PoisonedAlign. Future work includes developing more effective and robust defenses against PoisonedAlign. Further research may also investigate the trade-off between an attack’s potency and its generalization to craft even subtler and more diverse poisoning strategies. Finally, extending these poisoning techniques and corresponding defenses to the multi-modal domain represents a crucial step toward securing next-generation AI systems.
Acknowledgements.
We thank the reviewers for constructive feedback. This research was partially supported by NSF grant No. 2414406, 2131859, 2125977, 2112562, 1937787, and 2450935.